y=3x-2
y=x+2
Thanks! This is one question not two by the way~
Thanks!

Answer by checkley77(12844)

y=3x-2
y=x+2
3x-2=x+2
3x-x=2+2
2x=4
x=4/2
x=2 ans.
y=3*2-2
y=6-2
y=4 ans.
Proof:
4=2+2
4=4